OpenStudy (anonymous):

solve by using the quadratic formula 5x^2+2x=3

OpenStudy (anonymous):

first we must put all terms on the left side so itbecomes \(5x^2 + 2x - 3 = 0\) now we can use the quadratic formula... quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] so if we substitute... \[x = \frac{-2 \pm \sqrt{2^2 - 4(5)(-3)}}{2(5)}\] \[x = \frac{ -2 \pm \sqrt{4 +60}}{10}\] \[x = \frac{ -2 \pm \sqrt{64}}{10}\]

OpenStudy (anonymous):

\[x = \frac{-2 \pm 8}{10}\]

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

\(x = \frac{-2 \pm 8}{10}\) means that \(x = \frac{-2 + 8}{10}\) and \(x = \frac{-2 - 8}{10}\) therefore when \(x = \frac{-2 + 8}{10}\) the answer is \(x = \frac{-6}{10} = -\frac{3}{5}\) when \(x = \frac{-2 - 8}{10}\) the answer becomes \(x = \frac{-10}{10} = -1\)

OpenStudy (anonymous):

Thank you.